OpenStudy (anonymous):

Using complete sentences, explain which method you would use to solve the following system of equations and why. In your answer, include the solution to one of the variables and how you found it using the method you chose. x – 5y + 2z = 0 x + 4y – z = 12 2x – y + 3z = 10

OpenStudy (anonymous):

x - 5y + 2z = 0 x + 4y - z = 12 2x - y + 3z = 10 (x + x - 5y + 4y + 2z - z) = (0 + 12) 2x - y + z = 12 2x - y + z = 12 2x - y + 3z = 10 2x - y + z - 2x + y - 3z = 12 - 10 -2z = 2 z = -1 x - 5y + 2z = 0 x - 5y - 2 = 0 x + 4y - z = 12 x + 4y - 1 = 12 x = 5y + 2 x = 13 - 4y 5y + 2 = 13 - 4y 9y = 11 y = 11/9 x = 5y + 2 x = 55/9 + 2 x = 73/9 (73/9 , 11/9 , -1)

OpenStudy (anonymous):

Did you use elimination or substitution?

OpenStudy (anonymous):

I use the Elimination and substitution methods. You can use matrices

OpenStudy (anonymous):

But I'm not good at those

OpenStudy (anonymous):

Lol well thank you either way :) You're better than I am.